OpenStudy (anonymous):

Determine all points of intersection y=cosX and y=sinX in the first quadrant

OpenStudy (anonymous):

o my i wish i know..sorry i don't know

OpenStudy (arindameducationusc):

easy

OpenStudy (arindameducationusc):

sin45=cos45

OpenStudy (arindameducationusc):

hope that works

OpenStudy (anonymous):

Yes I know it's 45(pi/4) but is there any math involved? I found the answer by looking at a graph...

OpenStudy (arindameducationusc):

in terms of triangle... well try imagining with random values in which sin and cos value merge....

OpenStudy (aakashsudhakar):

Well, the more accurate answer would be:\[\frac{ \pi }{ 4 }+2n \pi\]This is because every 2*pi you go out in the positive direction, you'll have another intersection of the cosine and sine graphs in the first quadrant. This is a property generally held true for any intersection relationships between sinusoidal functions.

OpenStudy (arindameducationusc):

@AakashSudhakar How will you explain why the pi/4 came? We know its pi/4.

OpenStudy (usukidoll):

maybe it's from the 45 degree angle special triangle?

OpenStudy (usukidoll):